20x6-7x5-6x4-500x4+175x3+150x2 Factoring this equation?
Answer by richwmiller(17219)

20x^6-7x^5-6x^4-500x^4+175x^3+150x^2
First take out x^2
x^2*(20x^4-7x^3-6x^2-500x^2+175x+150)
This conveniently arranged into two parts- notice the x^2 twice -6x^2-500x^2
20x^4-7x^3-6x^2=x^2(4x-3)(5x+2)
-500x^2+175x+150=-25(4x-3)(5x+2)
so now we have
x^2*(4x-3)(5x+2)(x^2-25)
and finally
x^2*(x-5)*(x+5)*(4x-3)*(5x+2)